### v1

Abstract
Explainable artificial intelligence (XAI) has become increasingly important in decision-critical domains such as healthcare, finance, and law. Counterfactual (CF) explanations, a key approach in XAI, provide users with actionable insights by suggesting minimal modifications to input features that lead to different model outcomes. Despite significant advancements, existing CF generation methods often struggle to balance proximity, diversity, and robustness, limiting their real-world applicability. A widely adopted framework, Diverse Counterfactual Explanations (DiCE), emphasizes diversity but lacks robustness, making CF explanations sensitive to perturbations and domain constraints. To address these challenges, we introduce DiCE-Extended, an enhanced CF explanation framework that integrates multi-objective optimization techniques to improve robustness while maintaining interpretability. Our approach introduces a novel robustness metric based on the Dice-Sørensen coefficient, ensuring stability under small input variations. Additionally, we refine CF generation using weighted loss components ($\lambda_{p}$, $\lambda_{d}$, $\lambda_{r}$) to balance proximity, diversity, and robustness. We empirically validate DiCE-Extended on benchmark datasets (COMPAS, Lending Club, German Credit, Adult Income) across multiple ML backends (Scikit-learn, PyTorch, TensorFlow). Results demonstrate improved CF validity, stability, and alignment with decision boundaries compared to standard DiCE-generated explanations. Our findings highlight the potential of DiCE-Extended in generating more reliable and interpretable CFs for high-stakes applications. Future work will explore adaptive optimization techniques and domain-specific constraints to further enhance CF generation in real-world scenarios.
Keywords: Counterfactual Explanations · Machine Learning · Interpretability · XAI.
1 Introduction
The field of explainable artificial intelligence (XAI) has gained significant attention in recent years, particularly in domains where transparency and accountability are critical. One key approach within XAI is counterfactual (CF) explanations, which provide individuals with actionable insights by suggesting modifications to input variables that would lead to a different model prediction. However, existing CF generation methods often struggle with balancing proximity, diversity, and robustness [33], [27], [15]. The need for precise, interpretable, and adaptable CFs is particularly evident in high-stakes decision-making contexts such as healthcare, finance, and education.
A widely used framework in this domain is DiCE (Diverse Counterfactual Explanations) [33], which emphasizes diversity by generating multiple alternative CFs. For example, in a financial setting, DiCE might suggest different strategies for improving loan approval chances, such as increasing income or reducing debt. Despite its advantages, DiCE exhibits certain limitations, particularly in ensuring robustness under perturbations and incorporating causal relationships into CF generation [40], [20]. These shortcomings can lead to explanations that, while diverse, may lack feasibility or stability when applied in real-world settings.
Recent advancements have sought to improve CF generation by integrating causal reasoning, optimization strategies, and reinforcement learning [29], [21]. Techniques such as Sparse CounterGAN leverage generative adversarial networks (GANs) to ensure that generated CFs align with real-world constraints [40]. Other approaches, such as Multi-Objective CFs, employ multi-objective optimization to balance trade-offs between proximity, diversity, and robustness [8]. Additionally, reinforcement learning (RL)-based frameworks, such as Soft Actor-Critic RL for CFs, utilize actor-critic architectures to optimize CF generation under dynamic constraints [13]. These methodologies demonstrate promising improvements, yet gaps remain in ensuring that CF explanations are both actionable and robust in practical applications [14], [37].
To address these challenges, we propose DiCE-Extended, an advanced framework designed to enhance the robustness of CF generation. Our contributions are as follows:
  1. 1.
    Robustness Enhancement: We integrate the Dice-Sørensen coefficient to measure CF stability under input perturbations [35].
  2. 2.
    Multi-Objective Optimization: We refine CF generation by integrating weighted parameters ($\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$) to optimize proximity, diversity, and robustness [15], [22]. These components are structured as:
    • Proximity: Ensuring that suggested changes are realistic and feasible.
    • Diversity: Providing multiple viable pathways to achieve the desired outcome.
    • Robustness: Maintaining stability in CF explanations under minor variations.
  3. 3.
    Empirical Machine Learning (ML) Validation: We evaluate DiCE-Extended on benchmark datasets (COMPAS, Lending Club, German Credit, Adult Income), demonstrating controllable performance over the original DiCE framework in fidelity, stability, and practical applicability [33], [40], [23].
The remainder of this paper is structured as follows: Section 2 discusses the theoretical foundations of CF generation and its associated challenges. Section 3 presents the methodology behind DiCE-Extended, including its optimization framework and robustness integration. Section 4 details experimental evaluations and performance comparisons. Finally, Section 5 concludes the study and outlines future research directions.
2 Background
The increasing adoption of ML in high-stakes domains such as healthcare, finance, and criminal justice necessitates transparent and interpretable models. However, complex models often function as “black boxes”, limiting trust in their predictions [18], [17]. XAI addresses this challenge through interpretability techniques like SHAP and LIME, which provide feature-based explanations but sometimes compromise fidelity [2], [28], [39]. Feature importance methods, including permutation importance, improve transparency by mitigating biases in variable relevance [1]. Additionally, tree-based models benefit from game-theoretic approaches for enhanced local and global interpretability [32]. Despite these advances, ensuring actionable and domain-relevant explanations remains a challenge, requiring integration of causal reasoning and balancing trade-offs between simplicity and accuracy [7], [38].
2.1 Approaches to Counterfactual Explanations
CF explanations provide “what-if” scenarios by suggesting minimal modifications to input features that alter a model’s prediction. Traditional CF explanations methods focus on perturbation vectors but often overlook feature dependencies and practical feasibility. Decision-tree-based techniques like CF Explanation Trees [25] improve interpretability, while Ordered CF explanations [26] prioritize feature modifications based on asymmetric dependencies. Distribution-aware methods such as DiCE enhance plausibility by incorporating feature correlations [24]. Cutting-edge approaches such as ReLAX leverage deep reinforcement learning for efficient and scalable counterfactual generation [5], [6], while interactive platforms like ViCE enhance user engagement through intuitive visualization tools [19]. However, despite these advancements, CF explanations remain susceptible to adversarial manipulation, highlighting the need for robust mechanisms that ensure reliability, fairness, and trust [4], [30].
2.2 Ensuring Robustness in Counterfactual Generation
Ensuring robustness in CF explanations is crucial for their reliability and practical application. Bansal et al. [3] demonstrated that robust training enhances the consistency of explanations, reducing sensitivity to model variations. Similarly, Datta et al. [9] highlighted that mitigating adversarial vulnerabilities strengthens the interpretability of CF explanations, making them more resistant to manipulation. To further enhance robustness, Huai et al. [huai2022automating] introduced an automated concept-based framework that provides certified robustness guarantees, minimizing the need for human intervention. Virgolin and Fracaros [36] proposed incorporating robustness constraints into CF generation, ensuring reliability even under input perturbations. Additionally, Bayesian frameworks for uncertainty quantification, as explored by Slack et al. [34], improve the stability of CF explanations by accounting for uncertainty in predictions. Finally, Dombrowski et al. [11] emphasized the importance of generating plausible CF explanations, which enhances fairness and ensures that suggested changes remain feasible within real-world constraints.
2.3 Challenges and Open Problems
Despite progress, CFEs face several persistent challenges, including robustness, actionability, scalability, and fairness, as summarized in Table 1.
Table 1: Summary of Challenges in CF Explanations.
Challenge Description and References
Robustness to Model Changes Ensuring CF explanations validity after model updates [27].
Balancing Objectives Optimizing sparsity, proximity, and diversity in CF explanations [8], [29].
Actionability Generating actionable CF explanations that adhere to real-world constraints [33].
Scalability Reducing computational costs for high-dimensional data [23], [16].
Framework Integration Embedding CF explanations into user-centric interpretability tools [12].
Ethical Implications Avoiding biases and ensuring fairness in CF explanations [10].
Evaluation Metrics Lack of standardized metrics for benchmarking CF explanations performance [15].
Complex Models Generating CF explanations for intricate architectures and deep learning (DL) models [31].
3 Methodology
3.1 Problem Definition and Input
This section introduces an approach that applies optimization strategies to improve interpretability, robustness, and feasibility in CF explanations while addressing existing challenges.
3.1.1 Objective:
Given an input instance $\mathbf{x}\in\mathbb{R}^{d}$ and a trained ML model $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$, the task is to generate a set of $k$ counterfactual examples ${\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{k}}$ that influence the model’s decision toward a specified target behavior while maintaining interpretability and actionability.
3.1.2 Input Components
The CF generation problem is formally defined by:
  • $\mathbf{x}$: The original input instance, represented as a $d$-dimensional feature vector.
  • $f$: A trained ML model, assumed to be differentiable, mapping inputs to real-valued predictions.
  • $k$: The number of CF examples to generate.
  • $\mathcal{C}$: The feasible space of CF candidates, restricted by domain constraints.
3.1.3 Output
A set of CF examples ${\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{k}}\subseteq\mathcal{C}$, each adhering to defined optimization criteria.
3.2 Hypotheses
The generation of CF explanations follows these hypotheses:
  1. 1.
    Feasibility Hypothesis: CF explanations close to the original input $\mathbf{x}$ in feature space are more actionable for users.
  2. 2.
    Diversity Hypothesis: A diverse set of CF explanations improves interpretability.
  3. 3.
    Robustness Hypothesis: CFs that remain stable under small perturbations to $\mathbf{x}$ are critical for ensuring reliability.
3.3 Mathematical Formalizations
To meet these hypotheses, the generation process is formulated as an optimization problem. Let $\mathbf{c}_{i}\in\mathcal{C}$ denote a CF candidate. The objective is to minimize the following proposed loss function:
$\begin{split}\mathcal{L}(\mathbf{c}_{1},\dots,\mathbf{c}_{k})=\underset{c_{1},...,c_{k}}{\text{arg min}}[{\mathcal{L}_{\text{y\_loss}}}+\lambda_{p}\mathcal{L}_{\text{Proximity}}-\\ \lambda_{d}\mathcal{L}_{\text{Diversity}}-\lambda_{r}\mathcal{L}_{\text{Robustness}}],\end{split}$ (1)
where $\lambda_{p}$, $\lambda_{d}$, and $\lambda_{r}$ are hyperparameters balancing proximity, diversity, and robustness, respectively.
3.3.1 Proximity Metric:
The proximity metric ensures minimal deviation from $\mathbf{x}$, computed as:
$\text{Proximity}=\frac{1}{k}\sum_{i=1}^{k}|\mathbf{c}_{i}-\mathbf{x}|_{p}.$ (2)
$\|\cdot\|_{p}$ represents a norm-based distance metric, such as:
  • $\ell_{1}$ norm: $\|\mathbf{c}_{i}-\mathbf{x}\|_{1}=\sum_{j=1}^{d}|c_{i,j}-x_{j}|$.
  • $\ell_{2}$ norm: $\|\mathbf{c}_{i}-\mathbf{x}\|_{2}=\sqrt{\sum_{j=1}^{d}(c_{i,j}-x_{j})^{2}}$.
Proximity can be incorporated into the objective function as a penalty term to minimize unnecessary deviations from the original input:
$\mathcal{L}_{\text{Proximity}}=\lambda_{p}\cdot\|\mathbf{c}_{i}-\mathbf{x}\|_{p},$ (3)
where $\lambda_{p}$ is a hyperparameter that controls the trade-off between proximity and other objectives, set to 0.5 as specified in the DiCE framework [33].
3.3.2 Diversity Metric:
Diversity among CF explanations is measured via a similarity kernel matrix $\mathbf{K}$:
$\text{Diversity}=\det(\mathbf{K}).$ (4)
where $\mathbf{K}$ is a similarity kernel matrix defined as:
$\mathbf{K}_{i,j}=\frac{1}{1+\text{dist}(\mathbf{c}_{i},\mathbf{c}_{j})},$ (5)
and $\text{dist}(\cdot,\cdot)$ is a distance metric, such as $\ell_{1}$ or $\ell_{2}$ norm, to measure the difference between CF explanations $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$.
Diversity can be explicitly incorporated into the loss function to encourage variability among CF explanations:
$\mathcal{L}_{\text{Diversity}}=-\lambda_{d}\cdot\det(\mathbf{K}),$ (6)
where $\lambda_{d}$ is a hyperparameter that regulates the trade-off between diversity and other objectives. A higher determinant value indicates a more diverse set of CF explanations. Following the DiCE framework [33], $\lambda_{d}=1.0$ is set to 1.0.
3.3.3 Robustness Metric:
Robustness is computed as:
$\text{Robustness}=\frac{2\cdot|\mathbf{c}\cap\mathbf{c}^{\prime}|}{|\mathbf{c}|+|\mathbf{c}^{\prime}|},$ (7)
where $\mathbf{c}$ is CF explanations generated for $\mathbf{x}$ and $\mathbf{c}^{\prime}$ is perturbed CF explanations as $\mathbf{c}^{\prime}=\mathbf{c}+\boldsymbol{\delta}$, and $\boldsymbol{\delta}$ represents a slight noise, respectively.
During the robustness loss computation, counterfactuals and perturbed counterfactuals are binarized to ensure discreteness, then the resulting binary vectors are sent to Dice-Sørensen coefficient function. Robustness can be enhanced by introducing a regularization term into the objective function:
[…]
where $\lambda_{r}$ is a hyperparameter, and $robustness\_loss$ measures the distance between CF explanations with Dice-Sørensen coefficient.
We performed a grid search for the robustness weight $\lambda_{r}$ within the range $[0,1]$. As shown in Figure 1, the robustness loss reaches its peak at $\lambda_{r}=0.4$, while the total loss continues to decrease beyond this value. Additionally, the variation in computed robustness across the grid is minimal, approximately $3\times 10^{-4}$. To maintain a balanced trade-off between proximity, diversity, and robustness, we selected $\lambda_{r}=0.4$ as a moderate value. This ensures that robustness remains a contributing factor to the overall loss without dominating the optimization process.
Figure 1: Impact of robustness weight ($\lambda_{r}$) on robustness loss and total loss.
3.3.4 Constraints:
CF explanations must adhere to the following constraints:
  • Feature Constraints: Each feature value must lie within its valid range, specified by $\mathcal{C}$.
  • Actionability Constraints: Certain features may be immutable (e.g., gender or age) and must not be altered in the CF.
3.4 Dataset Information
To ensure the robustness of our CF explanation framework, we evaluate it on four diverse datasets [33].
  • Adult-Income: Extracted from the UCI Machine Learning Repository, this dataset includes demographic and employment-related features. We use 8 key attributes, such as hours per week, education level, and occupation, to classify individuals earning above or below \$50,000 per year.
  • LendingClub: This dataset contains five years of loan records (2007-2011) from LendingClub. We select 8 essential features, including employment years, annual income, and credit history, to predict whether a borrower will repay their loan.
  • German-Credit: This dataset includes 20 financial and demographic attributes related to loan applicants. The task is to classify applicants as good or bad credit risks.
  • COMPAS: A dataset from ProPublica focused on recidivism prediction in the U.S. criminal justice system. After preprocessing, 5 key features are retained (e.g., age, prior offenses, and charge severity) to predict whether a defendant will reoffend within two years.
3.5 Experimental Setup
To systematically evaluate the performance of the proposed CF explanation framework, we conduct experiments on all four datasets using three distinct ML frameworks: Scikit-learn, PyTorch (PYT), and TensorFlow (TF2). The evaluation follows a structured approach, ensuring fair comparisons across different models and loss function components.
3.5.1 Model Architectures
The experimental setup includes two distinct model architectures, each designed to assess CF explanations from different learning paradigms:
Tree-based Model.
We utilize a Random Forest classifier implemented with the Scikit-learn library. The model operates as a non-parametric ensemble learning method, aggregating multiple decision trees to enhance predictive stability. The hyperparameter settings enable reproducibility across experiments, with 100 estimators, the Gini criterion, and no maximum depth.
Neural Network Model.
To explore CF explanations in DL environments, we implement a two-layer feedforward neural network in both PYT and TF2. The architecture consists of a fully connected hidden layer with ReLU activation, followed by an output layer utilizing a Sigmoid activation function:
$f(\mathbf{x})=\sigma(W_{2}\cdot\text{ReLU}(W_{1}\mathbf{x}+b_{1})+b_{2}),$ (9)
where $W_{1}\in\mathbb{R}^{n\times d}$ and $W_{2}\in\mathbb{R}^{1\times n}$ are the weight matrices where $n$ is number of features, $b_{1}\in\mathbb{R}^{n}$ and $b_{2}\in\mathbb{R}$ are bias terms, and $\sigma$ represents the Sigmoid activation function. This neural network is trained via backpropagation using a binary cross-entropy loss function. The chosen architecture ensures model flexibility while maintaining computational efficiency during CF generation.
3.5.2 Training Configuration
To ensure consistency in model evaluation, we adhere to a standardized training protocol across all frameworks. The training process incorporates the following hyperparameter configurations:
  • Learning Rate: Set to 0.001, ensuring stable gradient updates while mitigating convergence issues.
  • Epochs: Fixed at 10 iterations, balancing training duration and model generalization.
  • Data Partitioning: An 80%-20% train-test split is applied, maintaining class distribution integrity within each dataset.
  • Batch Processing: Mini-batch gradient descent is utilized, with a batch size of 16 for training and 4 for testing.
  • Optimization Strategy: The Adam optimizer is employed, leveraging adaptive moment estimation to dynamically adjust learning rates.
To prevent overfitting, early stopping mechanisms are integrated, monitoring validation loss to halt training if no improvement is observed. Additionally, model evaluation metrics, including accuracy and loss progression, are continuously logged to assess the stability of CF explanations.
4 Experiments and Results
4.1 Loss Progression in CF Explainers
We analyzed the evolution of different loss components—including robustness loss—during the CF generation process across four datasets. Figure 2 illustrates how the optimization process minimizes these losses over multiple iterations, ensuring stable and interpretable CF explanations.
Focusing on robustness loss metric, we observe an initial fluctuation in the early iterations, followed by a steady decline as the optimization progresses. By approximately the 50th iteration, robustness loss stabilizes, indicating that the generated CFs maintain consistency despite small perturbations in input features. This trend suggests that the optimization framework successfully enhances the resilience of CF explanations, making them more reliable for real-world applications. The convergence of robustness loss aligns with the overall loss minimization strategy, balancing proximity, diversity, and stability without compromising interpretability. Moreover, the stabilization of robustness loss demonstrates that after a certain number of iterations, further optimization does not significantly improve CF reliability. This finding highlights the efficiency of the framework in generating CFs within a limited number of steps while ensuring robustness in decision-making scenarios.
4.2 Performance Comparison: Original vs. DiCE vs. DiCE-Extended Scenarios
To evaluate the effectiveness of DiCE-Extended, we compare ML model performance across three conditions: (1) models trained and tested on the original datasets, (2) models trained on original datasets but tested on CF datasets generated by DiCE, and (3) models trained on original datasets but tested on CF datasets generated by DiCE-Extended. The comparison focuses on key classification metrics, including accuracy, F1-score, recall, precision, AUC, as well as generation time and validity scores of CF samples.
4.2.1 Key Observations and Findings
1. Original Model Performance (Baseline Evaluation):
Models trained and tested on the original datasets generally exhibit high accuracy, particularly in datasets like Adult Income and Lending Club, where scores exceed 80% accuracy in most cases (Table 2). However, performance varies across different backends (Scikit-learn, PYT, TF2) due to their respective architectures and optimization mechanisms.
2. DiCE CFs (Baseline CF Evaluation):
When models trained on the original datasets were tested on CF examples generated by DiCE, there was a notable performance degradation. While Scikit-learn models retained perfect accuracy across datasets, DL models in PYT and TF2 showed significant drops in accuracy and F1-score, particularly in high-dimensional datasets like German Credit (Table 3). This suggests that standard DiCE-generated CFs may not always align with the decision boundaries of DL models, impacting model validity.
3. DiCE-Extended CFs (Optimized CF Evaluation):
The introduction of DiCE-Extended resulted in more stable CF explanations, leading to improved validity and classification performance compared to the original DiCE framework (Table 4). Notably:
  • PYT models showed improved robustness, with higher F1-scores and AUC values compared to their performance with standard DiCE CFs.
  • TF2 models demonstrated increased validity, particularly in high-dimensional datasets like COMPAS-Recidivism, where DiCE-Extended CFs aligned better with the decision function.
  • Generation time increased for TF2-based models, reflecting the additional optimization complexity introduced by robustness constraints in DiCE-Extended.
Figure 2: Loss trends in DiCE-Extended across datasets, illustrating the optimization of class loss, proximity loss, diversity loss, and robustness loss during CF generation.
Table 2: Classification performance of ML models trained and evaluated on original datasets.
Dataset Backend Accuracy F1-score Recall Precision AUC
Adult-income Scikit-learn 0.82 0.56 0.48 0.68 0.70
PYT 0.83 0.77 0.76 0.78 0.89
TF2 0.83 0.61 0.55 0.69 0.74
Compas-recidivism Scikit-learn 0.57 0.58 0.60 0.57 0.57
PYT 0.66 0.66 0.66 0.66 0.72
TF2 0.66 0.65 0.64 0.66 0.66
German-credit Scikit-learn 0.72 0.41 0.33 0.54 0.61
PYT 0.79 0.73 0.72 0.76 0.81
TF2 0.76 0.53 0.45 0.64 0.67
Lending-club Scikit-learn 0.82 0.90 0.99 0.83 0.51
PYT 0.83 0.46 0.50 0.60 0.69
TF2 0.83 0.91 1.00 0.83 0.51
Table 3: Classification performance of ML models trained on original datasets and evaluated on CF datasets (DiCE).
Dataset Backend Accuracy F1-score Recall Precision AUC Time (s) Validity
Adult-income Scikit-learn 1.00 1.00 1.00 1.00 1.00 1.00 0.77
PYT 0.56 0.53 0.56 0.58 0.55 0.59 0.89
TF2 0.41 0.11 0.07 0.22 0.41 0.52 0.88
Compas-recidivism Scikit-learn 1.00 1.00 1.00 1.00 1.00 1.00 0.49
PYT 1.00 1.00 1.00 1.00 1.00 0.88 0.18
TF2 0.44 0.41 0.38 0.44 0.44 0.68 0.49
German-credit Scikit-learn 1.00 1.00 1.00 1.00 1.00 0.14 0.70
PYT 0.58 0.56 0.58 0.61 0.63 0.87 0.49
TF2 0.54 0.43 0.35 0.56 0.54 0.17 0.58
Lending-club Scikit-learn 1.00 1.00 1.00 1.00 1.00 0.67 0.67
PYT 0.49 0.33 0.25 0.50 0.50 0.18 0.48
TF2 0.72 0.77 0.96 0.65 0.72 0.85 0.58
Table 4: Classification performance of ML models trained on original datasets and evaluated on CF datasets (DiCE-Extended).
Dataset Backend Accuracy F1-score Recall Precision AUC Time (s) Validity
Adult-income Scikit-learn 1.00 1.00 1.00 1.00 1.00 0.20 0.89
PYT 0.47 0.39 0.47 0.43 0.46 0.61 0.97
TF2 0.51 0.29 0.20 0.54 0.52 0.27 0.99
Compas-recidivism Scikit-learn 1.00 1.00 1.00 1.00 1.00 0.11 0.66
PYT 0.85 0.85 0.86 0.87 0.97 0.08 0.58
TF2 0.38 0.43 0.46 0.40 0.38 0.10 0.90
German-credit Scikit-learn 1.00 1.00 1.00 1.00 1.00 0.49 0.45
PYT 0.61 0.61 0.62 0.63 0.69 0.10 0.48
TF2 0.55 0.42 0.33 0.59 0.55 0.64 0.99
Lending-club Scikit-learn 1.00 1.00 1.00 1.00 1.00 0.26 0.77
PYT 0.48 0.32 0.48 0.24 0.23 0.18 0.76
TF2 0.70 0.76 0.91 0.65 0.71 0.10 0.87
5 Conclusion
Our quantitative analysis highlights that DiCE-Extended improves the robustness, validity, and overall reliability of CF explanations across different ML backends. While DiCE-generated CFs introduce notable performance degradation, DiCE-Extended mitigates these issues, ensuring better alignment with decision boundaries and producing more actionable CFs. Future work will focus on optimizing multi-objective trade-offs, integrating domain-specific constraints, and improving computational efficiency, making DiCE-Extended a more practical and interpretable counterfactual generation framework for real-world applications.
Disclosure of Interests.
The authors have no competing interests to declare that are relevant to the content of this article.